# Boosting Semi-Supervised 2D Human Pose Estimation by Revisiting Data Augmentation and Consistency Training

### v2

@@ 4 Our Overall Framework: MultiAugs
$\mathcal{L}^{*}_{u,A}=\mathbb{E}_{\mathbf{H}_{h_{i},A}\in\mathcal{H}_{A}}\sum\nolimits^{n}_{i=1}||\mathbf{H}_{e,B}-\mathbf{H}_{h_{i},A}||^{2}$ (7)
$\mathcal{L}^{*}_{u,B}=\mathbb{E}_{\mathbf{H}_{h_{i},B}\in\mathcal{H}_{B}}\sum\nolimits^{n}_{i=1}||\mathbf{H}_{e,A}-\mathbf{H}_{h_{i},B}||^{2}~{}$
@@
where we swap positions of teacher signals $\mathbf{H}_{e,A}$ and $\mathbf{H}_{e,B}$ for realizing the cross training of networks $f_{B}$ and $f_{A}$. Following [58], we report the average accuracy of the final two well-trained and performance-approached models. Besides, $f_{A}$ and $f_{B}$ can have different structures as in [58, 23], where the large one often helps to distill a better small model, but not vice versa. We do not intend to explore this consensus anymore in this paper.
5 Experiments
5.1 Datasets and Metrics
@@
COCO. The dataset COCO [34] has 4 subsets: train-set (118K images), val-set (5K images), test-dev and test-challenge. It is a popular large-scale benchmark for human pose estimation, which contains over 150K annotated people. In addition, there are 123K wild unlabeled images (wild-set). We randomly selected 1K, 5K and 10K labeled data from train-set. In some experiments, unlabeled data came from the remaining images of train-set. In other experiments, we used the whole train-set as the labeled dataset and wild-set as the unlabeled dataset. The metric of mAP (Average AP over 10 OKS thresholds) is reported.
MPII and AI-Challenger. The dataset MPII [1] has 25K images and 40K person instances with 16 keypoints. The dataset AI-Challenger (AIC) [53] train-set has 210K images and 370K person instances with 14 keypoints. We use MPII as the labeled set, AIC as the unlabeled set. The metric of PCKh0.5 is reported.
5.2 Implementation Details
@@
For a fair comparison with prior works, we use SimpleBaseline [55] to estimate heatmaps and ResNet [20] and HRNet [49] as backbones. The input image size is set to $256\times 192$. We adopt the PyTorch 1.30 and 4 A100 GPUs with each batch size as 32 for training. The initial learning rate is 1e-3. When training on COCO with 10K labeled data, it decreases to 1e-4 and 1e-5 at epochs 70 and 90, respectively, with a total of 100 epochs. When using 1K or 5K labeled data, total epochs are reduced to 30 or 70, respectively. When training on the complete COCO or MPII+AIC, it drops to 1e-4 and 1e-5 at epochs 220 and 260, respectively, with a total of 300 epochs. When testing, we do not flip horizontally.
@@
For data augmentation settings, we keep the easy augmentation $T_{e}$ as $T_{A30}$ in all experiments. The hard augmentations set $\mathcal{\widehat{T}}^{4}\!=\!\{T_{JOCO},T_{JCCM},T_{JC},T_{JO}\}$ is chosen for the comparison C1. While, the less optimal augmentations set $\mathcal{\widehat{T}}^{2}\!=\!\{T_{JOCO},T_{JCCM}\}$ is chosen for comparisons C2, C3, C4 and C5 to balance the performance and training time. We present details of finding two novel superior hard augmentations $T_{JOCO}$ and $T_{JCCM}$, and selecting the optimal augmentation set $\mathcal{\widehat{T}}^{4}$ in ablation studies.
5.3 Performance Comparison
@@
Table 3: AP of different methods on COCO val-set when different numbers of labels are used. Backbone is ResNet18.
  Method Net. Num. 1K 5K 10K All
@@
Figure 8: Comparison between state-of-the-art SSHPE methods and our proposed MultiAugs on the COCO dataset.
Table 4: Results on the COCO val-set when labeled images from the train-set and unlabeled images from the wild-set.
  Method Backbone Networks AP AP.${}_{.5}$ AR AR.${}_{.5}$
@@
Ours (Single) ResNet50 Single 74.0 92.5 77.1 93.9
@@
Ours (Single) ResNet101 Single 75.4 93.4 78.4 94.1
Ours (Dual) ResNet101 Dual 76.3 93.5 79.2 94.4
@@
Ours (Single) HRNet-w48 Single 79.3 94.6 81.8 95.1
@@
Table 5: Comparison to the SOTA methods on the COCO test-dev. The COCO train-set is the labeled set and COCO wild-set is the unlabeled set. The person detection results are provided by SimpleBaseline and flipping strategy is used.
  Method Backbone Input Size GFLOPS #Params AP AR
@@
ViTPose [59] ViT-Large 256$\times$192 59.8 307.0 77.3
@@
Ours (Dual) (+DARK) HRNet-w48 384$\times$288 65.8 127.2 77.3 82.3
Table 6: The results on the val-set of the MPII dataset. HRNet is trained only on the MPII train-set. The “*” means using extra labeled dataset AIC. The “+” means applying the model ensemble.
@@
Table 7: Comparisons on the MPII test-set (PCKh0.5). We use HRNetW32 as the backbone. Input size is 256$\times$256. The MPII and AIC (w/o labels) are used for training.
@@
We mainly compare our MultiAugs with representative SSHPE methods PoseDual [58] and SSPCM [23] under five conditions:
C1: Firstly, we conduct experiments on the COCO train-set with 1K, 5K and 10K labeled data, and evaluate on the val-set. As shown in Tab. 5.3 and Fig. 5.3, our method brings substantial improvements under the same setting. For example, when using a Single-Network, our method exceeds both PoseCons and PoseDual significantly, and is close to the SSPCM based on three networks. When using a Dual-Network, our method exceeds previous SOTA results by 2.8 mAP, 1.3 mAP, and 1.1 mAP under 1K, 5K and 10K settings, respectively. Note that our method can bring greater gains with less labeled data (e.g., 1K images), which further explains its efficiency and superiority.
C2: Then, we conduct larger scale SSHPE experiments on the complete COCO dataset by using train-set as the labeled dataset and wild-set as the unlabeled dataset. As shown in Tab. 4, regardless of using any backbone, our method can always improve all supervised baseline results, and bring more gains than two compared SSHPE methods [58] and [23] with using dual networks. When using a single network, our method is still superior to PoseDual [58], and fairly close to SSPCM [23] based on triple networks.
C3: Besides, we also report our results using HRNet-w48 on the COCO test-dev in Tab. 5. Our method can slightly outperform the PoseDual but fall behind the best SSHPE. We attribute it to our fewer training epochs and less parameters which may lead to weaker generalization. Besides, we can observe that MultiAugs outperforms some burdensome transformer-based methods [64, 32, 65], which reveals the significance of rational utilization of unlabeled data.
@@
C4: Finally, we allocate the MPII train-set as a labeled set and the whole AIC as an unlabeled set. This is a more realistic setting where labeled and unlabeled images are from different datasets. Tab. 6 shows results on the MPII val-set. Our method outperforms both the supervised HRNet [49] and semi-supervised PoseDual [58] by a large margin under the same backbone. It is worth noting that our semi-supervised MultiAugs with applying the model ensemble can even approach the supervised HRNet with using extra labeled AIC.
@@
C5: Then, as shown in Tab. 7, our results on the MPII test-set surpass all those of previous fully supervised methods and two semi-supervised counterparts PoseDual and SSPCM. This further validates the effectiveness and superiority of our proposed method.
5.4 Ablation Studies
To further study our proposed MultiAugs, we conduct ablation experiments under the setting of COCO 1K with a total of 30 epochs. The learning rate drops 10$\times$ twice separately at epochs 20 and 25. All studied models take ResNet18 as the backbone, and follow the Single-Network framework.
Table 8: Best mAP results of different augmentations on COCO val-set.
  Augmentations Year mAP
$T_{RA}$ [11] CVPRW’2020 41.9
$T_{TA}$ [38] ICCV’2021 40.2
$T_{YOCO}$ [17] ($T_{RA}$) ICML’2022 42.5
$T_{YOCO}$ [17] ($T_{TA}$) ICML’2022 41.6
$T_{JC}$ [58] ICCV’2021 42.4
$T_{JO}$ [23] CVPR’2023 41.9
Ours ($T_{JCCM}$) 42.7
Ours ($T_{JOCO}$) 43.7
Figure 9: The convergence curves and best mAP results of different strong augmentations.
5.4.1 Comparing to Other Advanced Augmentations.

### v3

@@ IV-2 MultiAugs (Dual-Network)
$\mathcal{L}^{*}_{u,A}=\mathbb{E}_{\mathbf{H}_{h_{i},A}\in\mathcal{H}_{A}}\sum\nolimits^{n}_{i=1}||\mathbf{H}_{e,B}-\mathbf{H}_{h_{i},A}||^{2},$ (7)
$\mathcal{L}^{*}_{u,B}=\mathbb{E}_{\mathbf{H}_{h_{i},B}\in\mathcal{H}_{B}}\sum\nolimits^{n}_{i=1}||\mathbf{H}_{e,A}-\mathbf{H}_{h_{i},B}||^{2},~{}$
@@
where we swap positions of teacher signals $\mathbf{H}_{e,A}$ and $\mathbf{H}_{e,B}$ for realizing the cross training of networks $f_{B}$ and $f_{A}$. Following [17], we report the average accuracy of the final two well-trained and performance-approached models. Besides, $f_{A}$ and $f_{B}$ can have different structures as in [17, 20], where the large one often helps to distill a better small model, but not vice versa. We thus do not explore this consensus anymore.
V Experiments
V-A Datasets and Setups
V-A1 Benchmarks
@@
Our experiments are mainly conducted on several HPE datasets such as COCO [15], MPII [14], AI-Challenger [16], CEPDOF [67], WEPDTOF [68] and COCO-WholeBody Hands [69]. The first three benchmarks are mainly used to study 2D human keypoint under conventional undistorted cameras. Specifically, the dataset COCO [15] has 4 subsets: train-set (118K images), val-set (5K images), test-dev and test-challenge. It is a popular large-scale benchmark for human pose estimation, which contains over 150K annotated people with each represented by 17 keypoints. In addition, there are 123K wild unlabeled images (wild-set). We selected the first 1K, 5K and 10K samples from train-set as the labeled set. In some experiments, unlabeled data came from the remaining images of train-set. In other experiments, we used the whole train-set as the labeled dataset and wild-set as the unlabeled dataset. The metric of mAP (Average AP over 10 OKS thresholds) is reported. The dataset MPII [14] has 25K images and 40K person instances with 16 keypoints. The dataset AI-Challenger (AIC) [16] train-set has 210K images and 370K person instances with 14 keypoints. We use MPII as the labeled set, AIC as the unlabeled set. The metric of PCKh0.5 is reported.
To further verify the effectiveness of our method, we conducted experiments on an indoor overhead fisheye human keypoint dataset WEPDTOF-Pose which is based on CEPDOF [67] and WEPDTOF [68]. Following SSPCM [20], we used the complete WEPDTOF-Pose train-set (4,688 person instances) as the labeled dataset, and CEPDOF [67] with 11,878 person instances as the unlabeled dataset for experiment. The WEPDTOF-Pose test-set (1,179 person instances) is used as the evaluation set. The metric of mAP [15] is reported for comparing. More details of WEPDTOF-Pose dataset can be found in [20]. It should be noted that SSPCM does not open source the BKFisheye dataset included in WEPDTOF-Pose, so we cannot conduct corresponding experimental comparisons involving the BKFisheye with it. Besides, to further verify the wide applicability and superiority of MultiAugs, we also conducted experiments on the human hand keypoint detection task. This is very similar to the human body pose estimation. Specifically, we selected relevant human hand images and corresponding keypoint annotations from the COCO-WholeBody dataset [69] as the annotated dataset (where the number of hand keypoints is 21), and obtained a train-set and a val-set containing approximately 76K and 3.8K samples, respectively. In addition, we used the BPJDet detector [70] to extract approximately 118K samples from COCO wild-set as the unlabeled dataset. This hand-oriented dataset can serve as a supplement to other trials of the whole body.
TABLE III: Detailed description of all seven training settings.
  Setting Name Labeled Dataset Unlabeled Dataset Evaluation Dataset Domain
S1 COCO train-set (front $x\%$) COCO train-set (left $1-x\%$) COCO val-set body + normal
S2 COCO train-set COCO wild-set COCO val-set body + normal
S3 COCO train-set COCO wild-set COCO test-set body + normal
S4 MPII train-set AIC train-set MPII val-set body + normal
S5 MPII train-set AIC train-set MPII test-set body + normal
S6 WEPDTOF- Pose train-set CEPDOF WEPDTOF- Pose test-set body + fisheye
S7 COCO Whole- Body train-set COCO wild-set COCO Whole- Body val-set hand + normal
V-A2 Training Settings
First, we formulate five SSL training experimental settings (including S1, S2, S3, S4 and S5) based on the arrangements in various previous SSHPE methods [17, 20, 21]. These setting mainly involve datasets COCO, MPII and AIC containing conventional human body keypoints. Then, we employ the sixth setting (S6) based on the dataset WEPDTOF-Pose, which mainly includes the images of human body keypoints with obvious distortion under the fisheye camera. Finally, we newly add the seventh setting (S7) focusing on the human hand keypoints detection task based on the dataset COCO-WholeBody Hands. The last two settings are domains that are different from mainstream 2D HPE tasks but have essential similarities, which is very suitable for us to demonstrate the universality of MultiAugs. We summarize the specifics of these settings in Tab. III.
V-A3 Implementation Details
@@
For a fair comparison with prior works, we use SimpleBaseline [55] to estimate heatmaps and ResNet [71] and HRNet [56] as backbones. The input image size is set to $256\times 192$. We adopt the PyTorch 1.30 and 4 A100 GPUs with each batch size as 32 for training. We use the Adam optimizer to train these models. The initial learning rate is 1e-3. When training on COCO with 10K labeled data, it decreases to 1e-4 and 1e-5 at epochs 70 and 90, respectively, with a total of 100 epochs. When using 1K or 5K labeled data, total epochs are reduced to 30 or 70, respectively. When training on the complete COCO or MPII+AIC, it drops to 1e-4 and 1e-5 at epochs 220 and 260, respectively, with a total of 300 epochs. When training on WEPDTOF-Pose, it decreases to 1e-4 and 1e-5 at 140 epochs and 180 epochs, respectively, with a total of 200 epochs. When training on COCO-WholeBody, we repeat the similar setting as the COCO dataset.
@@
For data augmentation, we keep the easy augmentation $T_{e}$ as $T_{A30}$. In Sec. III-A3 and Sec. III-B2, we have presented details of finding two novel superior hard augmentations (e.g., $T_{JOCO}$ and $T_{JCCM}$) and recommending the optimal multi-path augmentation set $\mathcal{\widehat{T}}^{4}$ (please see the scheme $m_{12}$ in Tab. III-B2), repsectively. Therefore, the hard augmentations set $\mathcal{\widehat{T}}^{4}\!=\!\{T_{JOCO},T_{JCCM},T_{JC},T_{JO}\}$ is chosen for settings S1, S2, S3, S6 and S7. While, the less optimal augmentations set $\mathcal{\widehat{T}}^{2}\!=\!\{T_{JOCO},T_{JCCM}\}$ is chosen for settings S4 and S5 to balance the performance and training time.
TABLE IV: Hyper-parameters details of basic augmentations.
  Aug. Type Description
$T_{A30}$ easy random scale within range $[0.75,1.25]$, random rotation within range $(-30^{\circ},30^{\circ})$.
$T_{A60}$ hard random scale within range $[0.75,1.25]$, random rotation within range $(-60^{\circ},60^{\circ})$.
$T_{CO}$ hard random generation 5 zero-valued size $20\times 20$ patches.
$T_{CM}$ hard random cropping 2 $20\times 20$ patches from other images.
$T_{JC}$ hard random generation 5 zero value patches with size $20\times 20$ around predicted joints.
$T_{JO}$ hard random cropping 2 patches from other images with size $20\times 20$ around predicted joints.
V-A4 Hyper-parameters of Basic Augmentations
The hyper-parameters involved in each augmentation are indeed important. In order to make a fair comparison, we list the parameters of the basic augmentations used in this paper in Tab. IV, so that readers can quickly and clearly know these details. Particularly, each basic augmentation we selected is derived from various compared methods without additional fine-tuning. For example, the parameters of Joint Cutout are the same as those in PoseDual [17] which used JC5, and the parameters of Joint Cut-Occlude are the same as those in SSPCM [20] which used JO2. Please refer Fig. 3 for some visualization results after applying augmentations.
V-B Performance Comparison
We mainly compare with representative SSHPE methods including PoseDual [17], SSPCM [20] and Pseudo-HMs [21] under various settings. Note that Pseudo-HMs does not follow the same setup as PoseDual [17] and SSPCM [20], and its code has not been released. We have tried our best to list partial comparable data in some tables to maintain completeness.
S1: Firstly, we conduct experiments on the COCO train-set with 1K, 5K and 10K labeled data, and evaluate on the val-set. As shown in Tab. V, our method brings substantial improvements under the same setting. For example, when using a Single-Network, our method exceeds both PoseCons and PoseDual significantly, and is close to the SSPCM based on three networks. When using a Dual-Network, our method exceeds previous SOTA results by 2.8 mAP, 1.3 mAP, and 1.1 mAP under 1K, 5K and 10K settings, respectively. Note that our method can bring greater gains with less labeled data (e.g., 1K images), which further explains its efficiency and superiority. In addition, we consider that reporting the comparison results based on the more common ResNet50 is more convincing than ResNet18. Therefore, we replace the backbone in Tab. V with ResNet50 and re-conduct the comparative experiments. The results are shown in Tab. VI. Similar to using ResNet18, our method can still achieve a clear advantage. When using a single network, our method outperforms PoseCons and Posedual, while being comparable to SSPCM. And our dual-network based approach achieves significant advantages.
@@ V-B Performance Comparison
TABLE V: AP of different methods on COCO val-set when different numbers of labels are used. The backbone of all methods is ResNet18. The best results are in bold.
  Method Nets Labeled Samples
1K 5K 10K All
@@
TABLE VI: AP of different methods on COCO val-set when different numbers of labels are used. The backbone of all methods is ResNet50. The best results are in bold.
  Method Nets Labeled Samples
1K 5K 10K All
Supervised [55] 1 34.8 50.6 56.4 70.9
PoseCons [17] 1 43.1 57.2 61.8
PoseDual [17] 2 48.2 61.1 65.0
SSPCM [20] 3 49.8 61.8 65.5
Ours (Single) 1 49.3 61.4 65.2
Ours (Dual) 2 51.7 62.9 66.3
S2: Then, we conduct larger scale SSHPE experiments on the complete COCO dataset by using train-set as the labeled dataset and wild-set as the unlabeled dataset. As shown in Tab. VII, regardless of using any backbone, our method can always improve all supervised baseline results, and bring more gains than two compared SSHPE methods [17] and [20] with using dual networks. When using a single network, our method is still superior to PoseDual [17], and fairly close to SSPCM [20] based on triple networks.
S3: We also report results using HRNet-w48 on the COCO test-dev in Tab. VIII. The upper, middle and lower panels show CNN-based supervised, Transformers-based supervised, and SSHPE methods, respectively. Our method can slightly outperform the PoseDual but fall behind the best SSPCM. We attribute it to our fewer training epochs (300 vs. 400) and less parameters (2 Nets vs. 3 Nets) which may lead to weaker generalization. Besides, we can observe that MultiAugs outperforms some burdensome transformer-based methods [72, 73, 74], which reveals the significance of rational utilization of unlabeled data and advanced SSL techniques.
TABLE VII: Results on COCO val-set with labeled train-set and unlabeled wild-set for SSL training. The best results are in bold.
  Method Backbone Nets AP AP.5 AR AR.5
@@
Pseudo-HMs [21] ResNet50 Dual 74.1
Ours (Single) ResNet50 Single 74.4 93.6 77.4 94.0
@@
Pseudo-HMs [21] ResNet101 Dual 75.7
Ours (Single) ResNet101 Single 75.8 93.5 78.8 94.4
Ours (Dual) ResNet101 Dual 76.4 93.6 79.3 94.7
@@
Pseudo-HMs [21] HRNet-w48 Dual 79.4
Ours (Single) HRNet-w48 Single 79.3 94.6 81.9 95.1
@@
TABLE VIII: Comparison to SOTA methods on the COCO test-dev. Person detection results are provided by SimpleBaseline [55] and flipping strategy is used. The best and second best results are in bold and underlined, respectively.
  Method Backbone Input Size Gflops Params AP AR
@@
ViTPose [11] ViT-Large 256$\times$192 59.8 307.0 77.3 82.4
@@
Ours (Dual) (+DARK) HRNet-w48 384$\times$288 65.8 127.2 77.4 82.3
TABLE IX: Results on val-set of MPII. HRNet is trained only on MPII train-set. The “*” means using extra labeled dataset AIC. The “+” means applying the model ensemble. Best results are in bold.
@@
TABLE X: Results on the val-set of MPII dataset. The used backbone is ResNet18. The best results are in bold.
  Method Hea Sho Elb Wri Hip Kne Ank Total
Supervised [55] 89.6 84.8 72.0 58.4 57.8 49.4 41.2 65.3
PoseCons [17] 92.7 87.6 74.5 67.9 72.3 64.2 59.4 75.2
PoseDual [17] 93.3 88.4 75.0 67.3 72.6 65.3 59.7 75.6
SSPCM [20] 93.5 90.6 80.2 71.3 75.9 68.9 62.3 78.3
Ours (Single) 94.1 91.1 80.5 72.2 76.3 69.2 62.8 79.1
Ours (Dual) 94.7 92.4 81.2 73.3 76.8 70.6 63.9 79.7
TABLE XI: Comparisons on test-set of MPII. We use HRNet-w32 as the backbone. The input image size is 256$\times$256. The MPII (with labels) and AIC (without labels) are used for SSL training.
@@
S4: For the MPII dataset, we allocate its train-set as the labeled set and whole AIC as the unlabeled set. This is a more realistic setting where labeled and unlabeled images are from different datasets. The Tab. IX shows results reported on the MPII val-set. Our method outperforms both the fully supervised HRNet [56] and semi-supervised PoseDual [17] by a large margin with using the same backbone. It is worth noting that our semi-supervised MultiAugs with applying the model ensemble can even approach the supervised HRNet with using extra labeled AIC. Moreover, we deem that conclusions drawn from testing on small-scale data may not necessarily be generalized to other datasets. Therefore, we repeat the comparison in setting S1 and Tab. V by replacing the dataset from COCO into MPII. Specifically, we conduct experiments using the first 1K samples as labeled data and the left 39K samples as unlabeled data in MPII. The validation set of MPII is used to evaluate. The final comparison results are shown in Tab. X. Not surprisingly, our method still maintains a clear lead in performance, both in terms of overall accuracy and the specific accuracy of each joint. These experiments once again demonstrate that our method is indeed universally effective and superior across different datasets.
@@
S5: Finally, in Tab. XI, our results on MPII test-set surpass all those of previous fully supervised methods and two semi-supervised counterparts PoseDual and SSPCM. This further validates the effectiveness and superiority of our method.
TABLE XII: Comparison of our method to SOTA methods on the dataset WEPDTOF-Pose collected by the fisheye camera.
  Method Nets AP AR
Supervised [55] 1 49.5 53.4
PoseCons [17] 1 54.6 58.1
PoseDual [17] 2 55.1 59.0
SSPCM [20] 3 55.6 60.0
Ours (Single) 1 56.5 60.6
Ours (Dual) 2 57.1 61.3
S6: For dataset WEPDTOF-Pose, given the particularity of the fisheye data, the random rotation range used in all hard data augmentations is (-90, 90). It means that $T_{A60}$ is changed into $T_{A90}$. As shown in Tab. XII, our method always achieves the best AP and AR results whether using a single network or a dual network structure, surpassing the previous SOTA method SSPCM using a triple network. These experiments in the fisheye domain once again demonstrate the superiority and universality of our proposed MultiAugs.
TABLE XIII: AP of different methods on COCO-WholeBody Hands val-set when different numbers of labels are used.
  Method Nets Labeled Samples
1K 5K 10K
Supervised [55] 1 33.4 38.9 41.9
PoseCons [17] 1 52.1 57.4 59.4
PoseDual [17] 2 55.9 60.1 62.0
SSPCM [20] 3 58.8 62.4 64.5
Ours (Single) 1 56.3 60.9 64.1
Ours (Dual) 2 62.8 65.5 67.4
TABLE XIV: Results on the COCO-WholeBody Hands val-set with using unlabeled COCO wild-set for SSL training.
  Method Backbone Nets AP AR
Supervised [55] ResNet50 Single 62.1 74.9
PoseDual [17] ResNet50 Dual 65.9 78.3
SSPCM [20] ResNet50 Triple 66.3 78.8
Ours (Single) ResNet50 Single 66.8 79.2
Ours (Dual) ResNet50 Dual 67.4 79.7
Supervised [55] ResNet101 Single 64.5 76.8
PoseDual [17] ResNet101 Dual 68.0 80.2
SSPCM [20] ResNet101 Triple 69.1 81.1
Ours (Single) ResNet101 Single 69.9 81.7
Ours (Dual) ResNet101 Dual 71.8 83.3
S7: This setting is for training on COCO-WholeBody Hands. Firstly, we still take ResNet18 as the backbone as in the setup S1, and then conduct comparative experiments using the methods SimpleBaseline, PoseCons, PoseDual, SSPCM, and the proposed MutliAugs. The final conclusion is still impressive. As shown in Tab. XIII, when using a single-network structure, our method is still significantly better than the PoseCons using a single network or the PoseDual using two networks. When we use a dual-network structure, our method is significantly better than the SSPCM using a triple-network. These results and the trends are consistent with HPE performance on the whole body. Secondly, we follow the setup S2, and conduct experiments on COCO-WholeBody Hands by using the annotated train-set as the labeled set and the COCO wild-set as the unlabeled set. Without losing generality, we choose to perform various experiments under backbones ResNet50 and ResNet101. The quantitative results on the val-set are summarized in the Tab. XIV. From these results, we can find that our method still has an undoubted advantage, which is consistent with the phenomenon shown in other tests. In a word, the above experimental results show that the advanced augmentation combination and multi-path consistency loss strategy we proposed is indeed sustainable, effective and easy to promote. For example, for keypoint detection tasks, whether it is human body or hand, MultiAugs has reliable transferability and great potential in versatility. We expect that these experiments can further help to demonstrate the core contribution of this paper.
Qualitative Comparison: To make our advantages more intuitively demonstrated, we have presented qualitative visualization comparison results, mainly including the conventional human images from the COCO val-set and the fisheye camera images from the WEPDTOF-Pose dataset. We take use of the backbone ResNet18 for all compared methods to highlight the their performance differences. For models trained on COCO dataset, we use the label set with 10K samples for comparison. As shown in Fig. 8, pure supervised learning methods are usually prone to making mistakes or messing up, and other SSHPE methods do not perform well in some occlusion cases or edge keypoint detection. While, our method often obtains more accurate estimation results.
Figure 8: Qualitative results on COCO val-set (the first 6 examples) and WEPDTOF-Pose test-set (the last 3 examples). The predicted results of methods Supervised and SSPCM are directly fetched from the supplementary paper of SSPCM [20]. The details of the comparison between SSPCM and our predictions are highlighted in yellow circles for quick identification.
V-C Ablation Studies
For empirical studies in Fig. 2, III-A3, III-B2 and Tab. III-A3, III-B2, we conducted them follow the S1 using 1K labeled COCO images and the Single-Network framework. Below, with the same setting, we conducted more ablation experiments to further analyze our proposed MultiAugs.
TABLE XV: Best mAP results of different augmentations on COCO val-set.
  Augmentations mAP
$T_{RA}$ [26] (CVPRW’2020) 41.9
$T_{TA}$ [31] (ICCV’2021) 40.2
$T_{YOCO}$ [34]+$T_{RA}$ (ICML’2022) 42.5
$T_{YOCO}$ [34]+$T_{TA}$ (ICML’2022) 41.6
$T_{JC}$ [17] (ICCV’2021) 42.4
$T_{JO}$ [20] (CVPR’2023) 41.9
$T_{JCCM}$ (Ours) 42.7
$T_{JOCO}$ (Ours) 43.7
Figure 9: The corresponding convergence curves of different augmentations in Tab. V-C.
V-C1 Comparing to Other Specially Designed Advanced Augmentations